OpenStudy (anonymous):

Someone help please. What method do I use to solve this? p^m divided by p^n is equal to?

OpenStudy (anonymous):

\[\frac{ p^m }{ p^n }=?\] right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Exponent rules! p^m / (p^n) = p^(m-n)

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

Now I know the rule for it thank you :)